OpenStudy (anonymous):

2 r – 8 + 5 r = 10 + r solve equation and show work and steps explaining how you did it please?

OpenStudy (anonymous):

2r-8+5r=10+r 2r+5r-8=10+r 7r-8=10+r 7r-r-8=10 6r-8=10 6r=10+8 6r=18 r=18/6 r=3

OpenStudy (anonymous):

thank you :)